OpenStudy (romantic_ch3micals):

A = P (1+rt) And apprently I'm trying to get ready by itself so...can someone help?

OpenStudy (romantic_ch3micals):

Not ready I meant r

OpenStudy (chrishogan):

I can help you. :)

OpenStudy (romantic_ch3micals):

Yay! That's great..so what's the first step?

OpenStudy (chrishogan):

First you need to distribute. \[A=P(1+rt)\]\[A=P(rt+1)\]

OpenStudy (chrishogan):

If p were distributed what would you get.

OpenStudy (chrishogan):

\[A=Prt+P\]right?

OpenStudy (chrishogan):

Now we need to solve for r, so r needs to be divided on both sides.

OpenStudy (romantic_ch3micals):

Doesn't p need to be divided by both sides? No. Subtract then t divided to get r by itself?

OpenStudy (chrishogan):

Result: \[Result: \frac{ A }{ R }=Pt+P\]

OpenStudy (chrishogan):

Now multiply both sides by "A"\[Result: R=\frac{ A-P }{ Pt }\]

OpenStudy (chrishogan):

I hoped this helped you.

OpenStudy (romantic_ch3micals):

Ohhhhh ok I see it. Thank you